# Hybrid Quantum-inspired Resnet and Densenet for Pattern Recognition

### v6

@@ 4.2 Parameter learning of QRFNN and QDFNN
In the part 1 of the supplementary material, the framework of the MLP is also given. By comparison, it is found that the similarity of QRFNN, QDFNN and MLP lies in the feedforward layer-by-layer transfer of features. What’s more, the classical output layer with activation functions and biases further enhances the ability of our hybrid models for nonlinear problems [2]. It also makes sense to activate the quantum-inspired layers with varied activation functions. The advantage of the classical input and output layers of QRFNN and QDFNN is the convenience of changing the number of neurons. Hence, our hybrid models are capable of allowing inputs and outputs of different dimensions. In addition, on the basis of the chain rule [2], the gradients of loss functions to the parameter $\alpha_{r}^{l-1}$ and $\beta_{r}^{l-1}$ are given:
@@
where $loss$ and $loss^{\prime}$ denote the loss functions of QRFNN and QDFNN separately. The terms, $\frac{\partial O_{L+1-j}}{\partial O_{L-j}}$ and $\frac{\partial O_{L+1-j}^{\prime}}{\partial O_{L-j}^{\prime}}$ $(0\leq j\leq L-l$), are derivative matrices. And each element in each of the matrices involves additions and multiplications of many sine and cosine functions. Because the ranges of sine and cosine functions are both [-1,1], there are limitations of the absolute value of each element in each of the matrices, which makes the absolute value of each element not very large. As a result, the absolute values of $\frac{\partial(loss)}{\partial\alpha_{r}^{l-1}}$ and $\frac{\partial(loss)}{\partial\beta_{r}^{l-1}}$ will be not very large, either. On the other hand, a common problem at deep learning region, gradient explosion, is caused by very large values of the gradients of the loss functions to the parameters [2]. As a result, the parameters in the neural networks are unable to be updated and the neural networks are not capable of learning features of new data. However, our hybrid models may be able to prevent gradient explosion through this theoretical analysis. More details of the illustration of our hybrid models to prevent gradient explosion are in the part 4 of the supplementary material.
@@ 5 Numerical experiments
In this section, we validate the generalization power of our hybrid models with various datasets. We also illustrate the robustness of our hybrid models with various parameter attacks utilizing noise. The programming architecture we use is Pytorch. Two MLPs with different activation functions are used for comparison with QRFNN and QDFNN. The details of the MLPs are in part 1 and 10 of the supplementary material. HQNet and Two CNNs with different activation functions are used for comparison with QRCNN and QDCNN in generalization power and robustness test. In terms of HQNet proposed in 2024, it indicated higher accuracy and f1 score over AG-CNN and other advanced convolutional networks on large MRI image classification tasks [2, 27]. Its hybrid framework is similar with our QRCNN and QDCNN. Hence, we consider HQNet as the state-of-the-art hybrid models for comparison. Besides that, we also reduce the size of the original HQNet utilizing 2 qubits to apply to smaller-size image classification. SPDA module of the convolutional part are utilized in the original HQNet. Additionally, we use parallel spatial and channel attention with residual or dense connection in the convolutional part in our HQNet, which is exactly the same as that in QRCNN (see part 6 of the supplementary material for more details of the convolutional part). And we maintain the architectures of the random circuit of the feedforward part in our HQNet due to the distinguished performance of it [27]. We use ${\rm R}_{\rm x}$ and ${\rm CR}_{\rm x}$ gate to construct the random circuit in HQNet (see part 13 of the supplementary material) [27]. When it comes to the two classical CNNs, they are constituted by convolutional layers and fully connected layers with MLPs. And QRCNN is comprised of residually-connected convolutional layers and fully connected layers with QRFNN, and QDCNN is comprised of densely-connected convolutional layers and fully connected layers with QDFNN. The convolutional parts in CNNs and QRCNN are identical. And we set most of the hyperparameters including learning rate, the number of fully connected layers of our HQNet, CNNs, QDCNN and QRCNN the same to compare these models more accurately (see part 10 of the supplementary material). Moreover, an ablation study is also conducted to test the superiority of the residual and dense connections. The traditional quantum-inspired neural network without residual or dense connection is utilized for comparison in it. All the results in this chapter are testing results.
@@ 5.1 Generalization power test
Table 6: Accuracy on FASHIONMNIST with symmetrical noise
@@
mixed noise 0.9003 0.9011 0.9280
gaussian noise 0.9157 0.9194 0.9294
uniform noise 0.9165 0.9096 0.9380
Table 7: Accuracy on MNIST with symmetrical noise
@@
mixed noise 0.9472 0.9532 0.9680
gaussian noise 0.9506 0.9491 0.9794
uniform noise 0.9412 0.9405 0.9711
Table 8: Accuracy on FASHIONMNIST with unsymmetrical noise
@@
mixed noise 0.8957 0.8787 0.9270
gaussian noise 0.9003 0.9096 0.9201
uniform noise 0.9241 0.9190 0.9422
Table 9: Accuracy on MNIST with unsymmetrical noise
@@
mixed noise 0.9494 0.9305 0.9682
gaussian noise 0.9311 0.9435 0.9603
uniform noise 0.9487 0.9475 0.9622
@@
where True Positive $(TP)$ denotes the number of samples of a category truly predicted as positive if we regard this category as positive in a dataset and other categories in the dataset negative. True Negative $(TN)$ is the number of samples predicted as negative that is negative. False Positive $(FP)$ is the number of samples that are incorrectly predicted to be negative but are positive in fact. False Negative $(FN)$ represents the number of samples that are falsely predicted to be positive but are actually negative. The P-R curve plots precision against recall at different classification thresholds, while the ROC curve plots true positive rate against false positive rate at assorted classification thresholds. They are expected to be close to 1 [2]. Furthermore, we utilize iris data with three categories and MNIST, FASHIONMNIST, CIFAR100 image datasets with each image dataset comprised of four categories. The iris data is 4-dimensional, the size of MNIST and FASHIONMNIST is 28 $\times$ 28 $\times$ 1, while that of CIFAR dataset is 32 $\times$ 32 $\times$ 3. In terms of each dataset, we set 75% of it as training set and 25% of it as testing set. And the proportion of each category of data in each dataset is approximately the same. To emulate real-world scenarios more closely, the noises added are pure symmetrical and pure unsymmetrical gaussian noises, pure symmetrical and pure unsymmetrical uniform noises, and mixed noises with symmetrical and asymmetric distributions, a total of six cases [44, 45, 46]. The symmetrical noise means the noise data distribution which is symmetrical with respect to the y-axis in a Cartesian coordinate system (see supplementary material part 12) [45]. These noises used in the paper are significant in engineering control [44]. There are more types of asymmetric noise than symmetric noise [44]. Therefore, in the engineering field, the probability of asymmetric noise occurring is more than that of symmetric noise [44, 47]. Moreover, to add more uncertainties to simulate the situations of real world, the amount and the position of the noise added to the data are both random in each group of the experiment [45]. And in each group of generalization power experiment, the noisy dataset utilized for training and testing all the models is identical.
@@
Table 1 to Table 5 demonstrate the outcomes of FASHIONMNIST with pure gaussian symmetrical noise (more results of other datasets in supplementary materials part 7). The results include the P-R and ROC curve area and accuracy. They also incorporate variances and mean values of precision, recall and f1 score of the four categories, from category 0 to category 3. And the category Y (Y=0,1,2,3) means a category labeled as the number Y. The mean value is the average values of precision, recall and f1 score of all the samples in a certain category from the $100^{th}$ epoch to the $300^{th}$ epoch after convergence, which implies the learning ability of models for samples of different categories. The variance represents that of precision, recall and f1 score of all the samples in a certain category from the $100^{th}$ epoch to the $300^{th}$ epoch after convergence, which indicates the degree of fluctuation of the indicators, reflecting the stability of the models. As for Table 1, while the accuracy of QRCNN and the two CNNs is 92.84%, 92.99% and 93.10%, that of QDCNN is 93.16%, which is slightly higher than that of QRCNN and CNNs. However, the accuracy difference between these models is exceedingly small. And the differences of P-R and ROC curve area between the four models are very small as well. Table 2 shows the variances and mean values of recall, precision and f1 score of category 0 of noisy FASHIONMNIST with category 0 considered as positive and other categories regarded as negative. Table 3 shows the variances and mean values of recall, precision and f1 score of category 1 of noisy FASHIONMNIST with category 1 considered as positive and other categories regarded as negative. Table 4 shows the variances and mean values of recall, precision and f1 score of category 2 of noisy FASHIONMNIST with category 2 considered as positive. Table 5 shows the variances and mean values of recall, precision and f1 score of category 3 of noisy FASHIONMNIST with category 3 considered as positive. As for the Table 2 to Table 5, there are only very small discrepancy of the variances and mean values of the three metrics between the four models. And the outcomes in part 7 in the supplementary materials also indicate the exceedingly small discrepancy of those. Therefore, our hybrid models could not outperform pure classical models in all these indicators. They only perform approximately on par with classical CNNs on datasets with symmetrical noises on various metrics. What’s more, Table 6 and 7 demonstrate accuracy of HQNet, QRCNN as well as QDCNN on noisy FASHIONMNIST and MNIST respectively. The results point the slight superiority of QDCNN over QRCNN and HQNet. While QRCNN shows approxiamtely as high accuracy as the HQNet model, more significantly, the accuracy of QDCNN is about 2%-3% higher than HQNet and QRCNN models, which shows the ascendancy of the dense connection of the convolutional and quantum-inspired layers in QDCNN.
5.1.1 Test on datasets with unsymmetrical noises
@@
Table 8 and 9 show the accuracy of three models on noisy FASHIONMNIST and MNIST respectively. The outcomes also reveal the slightly more prominent performance of QDCNN with 2%-3% higher accuracy over HQNet and QRCNN. Other results of the comparison between our hybrid models and classical MLPs and CNNs are shown in the part 7 of the supplementary material. On the whole, the outcomes indicate that our hybrid models possess roughly the same level of generalization capability as MLPs or CNNs in recognizing data containing asymmetric noises.
@@ 5.2 Robustness test
Robustness is appraised by the same metrics in generalization power part. And in this part, we use the same datasets as in generalization power part. For QDFNN, QRFNN and the MLPs, in each group of robustness experiments, we randomly choose one layer and attack some parameters randomly in the layer utilizing the six noises in generalization power test. The layer indices and the number of parameters that are chosen for attacking in the layer of the three models are identical. In terms of QRCNN, QDCNN, CNNs and HQNet, since we hope to evaluate the robustness of the fully connected layers of these models, we randomly choose a fully-connected layers in the CNNs, a quantum-inspired layer in QRCNN and QDCNN, a quantum layer in HQNet for attacking. And we ensure the layer indices and the number of parameters attacked in the layer in CNNs is also exactly the same as that in HQNet and our QDCNN and QRCNN. In each group of robustness experiment, the noise used to attack different models is exactly the same. Nonetheless, the noise used for attacking and the number of parameters chosen in one layer for attacking between different groups of experiments are different. There are two common attacking forms: ① $\sigma(\theta+\epsilon)$ and ② $\sigma(\theta)+\epsilon$, where $\theta$ is the parameter and $\epsilon$ is the noise. $\sigma$ represents activation operations in pure classical models or sine and cosine functions in our hybrid models or HQNet[44, 48, 49, 50]. The test of robustness is separated into symmetrical noise attack with form ①,② and unsymmetrical noise attack with form ①,②. The accuracy value is the average accuracy value after the models reach convergence. The link to the program codes is on the last page of this paper.
5.2.1 Test under symmetrical noise attacking

### v7

@@ 4.2 Parameter learning of QRFNN and QDFNN
In the section 1 of the supplementary material, the framework of the MLP is also given. By comparison, it is found that the similarity of QRFNN, QDFNN and MLP lies in the feedforward layer-by-layer transfer of features. What’s more, the classical output layer with activation functions and biases further enhances the ability of our hybrid models for nonlinear problems [2]. It also makes sense to activate the quantum-inspired layers with assorted activation functions. The advantage of the classical input and output layers of QRFNN and QDFNN is the convenience of changing the number of neurons. Hence, our hybrid models are capable of allowing inputs and outputs of different dimensions. In addition, on the basis of the chain rule [2], the gradients of loss functions to the parameter $\alpha_{r}^{l-1}$ and $\beta_{r}^{l-1}$ are given:
@@
where $loss$ and $loss^{\prime}$ denote the loss functions of QRFNN and QDFNN separately. The terms, $\frac{\partial O_{L+1-j}}{\partial O_{L-j}}$ and $\frac{\partial O_{L+1-j}^{\prime}}{\partial O_{L-j}^{\prime}}$ $(0\leq j\leq L-l$), are derivative matrices. And each element in each of the matrices involves additions and multiplications of many sine and cosine functions. Because the ranges of sine and cosine functions are both [-1,1], there are limitations of the absolute value of each element in each of the matrices, which makes the absolute value of each element not very large. As a result, the absolute values of $\frac{\partial(loss)}{\partial\alpha_{r}^{l-1}}$ and $\frac{\partial(loss)}{\partial\beta_{r}^{l-1}}$ will be not very large, either. On the other hand, a common problem at deep learning region, gradient explosion, is caused by very large values of the gradients of the loss functions to the parameters [2]. As a result, the parameters in the neural networks are unable to be updated and the neural networks are not capable of learning features of new data. However, our hybrid models may be able to prevent gradient explosion through this theoretical analysis. More details of the illustration of our hybrid models to prevent gradient explosion are in the section 4 of the supplementary material.
@@ 5 Numerical experiments
In this section, we validate the generalization power of our hybrid models with various datasets. We also illustrate the robustness of our hybrid models with various parameter attacks utilizing noise. The programming architecture we use is Pytorch. All the experiments are from single runs. Two MLPs with different activation functions are used for comparison with QRFNN and QDFNN. The details of the MLPs are in section 1 and 10 of the supplementary material. HQNet and Two CNNs with different activation functions are used for comparison with QRCNN and QDCNN on generalization power and robustness test. Moreover, In terms of the original HQNet proposed in 2024, we contemplate it as the state-of-the-art hybrid models for comparison since its hybrid framework is similar with those of our QRCNN and QDCNN [2, 27]. Another reason lies in its more outstanding performance on several medical image classification tasks than numerous state-of-the-art classical convolutional networks such as DenseNet-121, MBTFCN, AG-CNN, etc. For instance, the experimental results on classification of BT-large-4c dataset indicate that while the accuracy of the advanced EfficientNetB0 and MobileNet is respectively 96.48% and 95.10%, that of the original HQNet is 97.55%. Simultaneously, the network model parameters and memory of the original HQNet is much less than that of EfficientNetB0 and MobileNet [27]. When it comes to the detailed architecture of the original HQNet, it is comprised of a hybrid residual module, a feature extraction and recombination module, a random quantum circuit module, and a classification module. And as for the HQNet in our experiments, we maintain the four modules but reduce the size of the random circuit model of the original HQNet from 4 qubits to 2 qubits to apply to smaller-size image classification. We use ${\rm R}_{\rm x}$ and ${\rm CR}_{\rm x}$ gate to construct the random circuit in our HQNet (see section 13 of the supplementary material) [27]. We also alter the number of the output neuron from 2 to 4 in the classification module for our multi-classification problem. What’s more, ref. [8] and [23] have shown neural networks with different residual mappings, all of which demonstrate outstanding performances. Hence, to further verify the ability of the residual and dense mapping our proposed hybrid models, we devise ResQNet1 and ResQNet2 by adjusting the residual mapping of QRCNN for comparison. The frameworks of ResQNet1, ResQNet2 and QRCNN are almost identical with the only discrepancy being the residual mapping in the quantum-inspired part. While one quantum-inspired layer constitute the residual mapping of QRCNN, that of ResQNet1 and ResQNet2 is comprised of all quantum-inspired layers and two quantum-inspired layers respectively. In addition, when it comes to the two classical CNNs, they consist of convolutional layers and fully connected layers with MLPs. And QRCNN is comprised of residually-connected convolutional layers and fully connected layers with QRFNN, and QDCNN is comprised of densely-connected convolutional layers and fully connected layers with QDFNN. The convolutional parts in CNNs and QRCNN are identical. And we set most of the hyperparameters including the number of fully connected layers of our HQNet, CNNs, QDCNN and QRCNN, learning rate, etc., the same to compare these models more accurately (see section 10 of the supplementary material for hyperparameter setting). Moreover, an ablation study is also conducted to test the superiority of the residual and dense connections. The traditional quantum-inspired neural network without residual or dense connection is utilized for comparison in it. All the results in this chapter are testing results.
@@ 5.1 Generalization power test
Table 6: Accuracy on FASHIONMNIST with three symmetrical noises
@@
mixed noise 0.9003$\pm$0.0105 0.9011$\pm$0.0102 0.9280$\pm$0.0094
gaussian noise 0.9157$\pm$0.0101 0.9194$\pm$0.0100 0.9294$\pm$0.0096
uniform noise 0.9165$\pm$0.0103 0.9096$\pm$0.0106 0.9380$\pm$0.0093
Table 7: Accuracy on MNIST with three symmetrical noises
@@
mixed noise 0.9472$\pm$0.0105 0.9532$\pm$0.0095 0.9680$\pm$0.0088
gaussian noise 0.9506$\pm$0.0095 0.9491$\pm$0.0097 0.9794$\pm$0.0082
uniform noise 0.9412$\pm$0.0106 0.9405$\pm$0.0107 0.9711$\pm$0.0083
Table 8: Accuracy on FASHIONMNIST with three asymmetrical noises
@@
mixed noise 0.8957$\pm$0.0103 0.8787$\pm$0.0111 0.9270$\pm$0.0096
gaussian noise 0.9003$\pm$0.0105 0.9096$\pm$0.0107 0.9201$\pm$0.0097
uniform noise 0.9241$\pm$0.0092 0.9190$\pm$0.0096 0.9422$\pm$0.0086
Table 9: Accuracy on MNIST with three asymmetrical noises
@@
mixed noise 0.9494$\pm$0.0098 0.9305$\pm$0.0101 0.9682$\pm$0.0086
gaussian noise 0.9311$\pm$0.0100 0.9435$\pm$0.0095 0.9603$\pm$0.0087
uniform noise 0.9487$\pm$0.0098 0.9475$\pm$0.0094 0.9622$\pm$0.0089
@@
where True Positive $(TP)$ denotes the number of samples of a category truly predicted as positive if we regard this category as positive in a dataset and other categories in the dataset negative. True Negative $(TN)$ is the number of samples predicted as negative that is negative. False Positive $(FP)$ is the number of samples that are incorrectly predicted to be negative but are positive in fact. False Negative $(FN)$ represents the number of samples that are falsely predicted to be positive but are actually negative. The P-R curve plots precision against recall at different classification thresholds, while the ROC curve plots true positive rate against false positive rate at assorted classification thresholds. They are expected to be close to 1 [2]. Furthermore, we utilize iris data with three categories and MNIST, FASHIONMNIST, CIFAR100 image datasets with each image dataset comprised of four categories. The iris data is 4-dimensional, the size of MNIST and FASHIONMNIST is 28 $\times$ 28 $\times$ 1, while that of CIFAR dataset is 32 $\times$ 32 $\times$ 3. In terms of each dataset, we set 75% of it as training set and 25% of it as testing set. And the proportion of each category of data in each dataset is approximately the same. To emulate real-world scenarios more closely, the noises added are pure symmetrical and pure asymmetrical gaussian noises, pure symmetrical and pure asymmetrical uniform noises, and mixed noises with symmetrical and asymmetrical distributions, a total of six cases [45, 46, 47]. The symmetrical noise means the noise data distribution which is symmetrical with respect to the y-axis in a Cartesian coordinate system (see supplementary material section 12) [46]. These noises used in the paper are significant in engineering control [45]. There are more types of asymmetrical noise than symmetrical noise [45]. Therefore, in the engineering field, the probability of asymmetrical noise occurring is more than that of symmetrical noise [45, 48]. Moreover, to add more uncertainties to simulate the situations of real world, the amount and the position of the noise added to the data are both random in each group of the experiment [46]. And in each group of generalization power experiment, the noisy dataset utilized for training and testing all the models is exactly identical. For instance, the four models are evaluated with identical FASHIONMNIST dataset with gaussian symmetrical noise in table 1.
5.1.1 Test on datasets with symmetrical noises
@@
Table 1 to Table 5 demonstrate the outcomes of FASHIONMNIST with pure gaussian symmetrical noise (more results of other datasets in supplementary materials section 7). The results include the P-R and ROC curve area and accuracy. They also incorporate variances and mean values of precision, recall and f1 score of the four categories, from category 0 to category 3. And the category Y (Y=0,1,2,3) means a category labeled as the number Y. The mean value is the average values of precision, recall and f1 score of all the samples in a certain category from the $100^{th}$ epoch to the $300^{th}$ epoch after convergence, which implies the learning ability of models for samples of different categories. The variance represents that of precision, recall and f1 score of all the samples in a certain category from the $100^{th}$ epoch to the $300^{th}$ epoch after convergence, which indicates the degree of fluctuation of the indicators, reflecting the stability of the models. As for Table 1, while the accuracy of QRCNN and the two CNNs is 92.84%$\pm$0.96, 92.99%$\pm$0.94and 93.10%$\pm$0.93, that of QDCNN is 93.16%$\pm$0.93, which is slightly higher than that of QRCNN and CNNs. However, the accuracy difference between these models is exceedingly small. And the differences of P-R and ROC curve area between the four models are very small as well. Table 2 shows the variances and mean values of recall, precision and f1 score of category 0 of noisy FASHIONMNIST with category 0 considered as positive and other categories regarded as negative. Table 3 shows the variances and mean values of recall, precision and f1 score of category 1 of noisy FASHIONMNIST with category 1 considered as positive and other categories regarded as negative. Table 4 shows the variances and mean values of recall, precision and f1 score of category 2 of noisy FASHIONMNIST with category 2 considered as positive. Table 5 shows the variances and mean values of recall, precision and f1 score of category 3 of noisy FASHIONMNIST with category 3 considered as positive. As for the Table 2 to Table 5, there are only very small discrepancy of the variances and mean values of the three metrics between the four models. And the outcomes in section 7 in the supplementary material also indicate the exceedingly small discrepancy of those. Therefore, our hybrid models could not outperform pure classical models in all these indicators. They only perform approximately on par with classical CNNs on datasets with symmetrical noises on various metrics. What’s more, Table 6 and 7 demonstrate accuracy of HQNet, QRCNN as well as QDCNN on noisy FASHIONMNIST and MNIST respectively. The results point the slight superiority of QDCNN over QRCNN and HQNet. While QRCNN shows approximately as high accuracy as the HQNet model, more significantly, the accuracy of QDCNN is about 2%-3% higher than HQNet and QRCNN models, which shows the ascendancy of the dense connection of the convolutional and quantum-inspired layers in QDCNN.
5.1.2 Test on datasets with asymmetrical noises
Table 10: Accuracy on CIFAR10 with three asymmetrical noises
noise type ResQNet1 ResQNet2 QRCNN QDCNN
uniform noise 78.04%$\pm$1.2977.49%$\pm$1.4776.82%$\pm$1.6279.33%$\pm$1.16%
gaussian noise 75.98%$\pm$1.7176.08%$\pm$1.6776.17%$\pm$1.6677.13%$\pm$1.64%
mixed noise 79.60%$\pm$1.2378.24%$\pm$1.3280.01%$\pm$1.2179.96%$\pm$1.22%
Table 11: Accuracy on MNIST with three asymmetrical noises
noise type ResQNet1 ResQNet2 QRCNN QDCNN
uniform noise 94.66%$\pm$0.7096.94%$\pm$0.5895.95%$\pm$0.6296.71%$\pm$0.59%
gaussian noise 93.86%$\pm$0.7895.42%$\pm$0.6695.05%$\pm$0.6996.22%$\pm$0.59%
mixed noise 93.23%$\pm$0.8593.97%$\pm$0.7994.01%$\pm$0.7694.73%$\pm$0.77%
@@ 5.1.2 Test on datasets with asymmetrical noises
Table 8 and 9 show the accuracy of HQNet, QRCNN and QDCNN on noisy FASHIONMNIST and MNIST respectively. The outcomes also reveal slightly more remarkable performance of QDCNN with 2%-3% higher accuracy over HQNet and QRCNN. Other results of the comparison between our hybrid models and classical MLPs and CNNs are shown in the section 7 of the supplementary material. On the whole, the outcomes indicate that our hybrid models possess roughly the same level of generalization capability as MLPs or CNNs in recognizing data containing asymmetrical noises.
What’s more, Table 10 and 11 showcase the accuracy of ResQNet1, ResQNet2 and our hybrid models on CIFAR10 and MNIST datasets with the three asymmetrical noises. The outcomes reveal a little bit more prominent performance of our hybrid models over ResQNet1 and ResQNet2, which demonstrates the slightly more prominent function of the residual mapping with one quantum-inspired layer.
@@ 5.2 Robustness test
Robustness is appraised by the same metrics in generalization power part. And in this part, we use the same datasets as in generalization power part. For QDFNN, QRFNN and the MLPs, in each group of robustness experiments, we randomly choose one layer and attack some parameters randomly in the layer utilizing the six noises in generalization power test. The layer indices and the number of parameters that are chosen for attacking in the layer of the three models are identical. In terms of QRCNN, QDCNN, CNNs, HQNet, ResQNet1 and ResQNet2, since we hope to evaluate the robustness of the fully-connected layers of these models, we randomly choose a fully-connected layers in the CNNs, a quantum-inspired layer in QRCNN, QDCNN, ResQNet1 and ResQNet2, a quantum layer in HQNet for attacking. And we ensure the layer indices and the number of parameters attacked in the layer in CNNs is also exactly the same as that in HQNet, QDCNN, QRCNN, ResQNet1 and ResQNet2. In each group of robustness experiment, the noise used to attack different models is identical. For example, in terms of each of the three models in Fig.(5a), one random quantum or one quantum-inspired layer with the same index is attacked by identical symmetrical mixed noise in form ①. Nonetheless, the noise used for attacking between different groups of experiments and the number of parameters chosen in one layer for attacking between various groups of experiments are different. There are two common attacking forms: ① $\sigma(\theta+\epsilon)$ and ② $\sigma(\theta)+\epsilon$, where $\theta$ is the parameter and $\epsilon$ is the noise. $\sigma$ represents activation operations in pure classical models or sine and cosine functions in our hybrid models or HQNet[45, 49, 50, 51]. The test of robustness is divided into symmetrical noise attack with form ①,② and asymmetrical noise attack with form ①,②. The accuracy value is the average one after the models reach convergence. The link to the program codes is on the last page of this paper.
5.2.1 Test under symmetrical noise attacks